G. Jorjadze

A. Razmadze Mathematical Institute, Georgian Academy of Sciences Tbilisi, Georgia

POISSON STRUCTURE OF WZNW THEORY

We investigate local and causal character of non-equal time Poisson brackets of SL(2, R) WZNW model and its coset theories. The quantum realisation is discussed.